If boys and girls are equally likely and groups of 400 births are randomly selected. Find the standard deviation of the numbers of girls in such groups of 400.
Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

This sounds like a binomial distribution problem.

Standard deviation:

sigma = sqrt(n*p*(1-p))

sigma = sqrt(400*0.5*(1-0.5))

sigma = 10

So the standard deviation is 10.

Answer by stanbon(75887)

If boys and girls are equally likely and groups of 400 births are randomly selected. Find the standard deviation of the numbers of girls in such groups of 400.
----
std = sqrt[npq] = sqrt[400(1/2)(1/2) = sqrt(100) = 10
===========================
Cheers,
Stan H.
===================